Find the electric force between two electrons separated by a distance of 10 cm.(The charge of an electron=-1.6x10^-19 C

1
Expert's answer
2022-03-04T08:30:25-0500

Answer

the electric force between two electrons separated by a distance of 10 cm is given by

"F=\frac{kqq'}{r^2}\\=\frac{9*10^9*1.6*1.6*10^{-38}}{0.1*0.1}\\=23.04*10^{-27}N"